Solve the equation (if possible) and check your soltuion.
4(x+3)-3=2(4-3x)-4

Answer by terbear1963(43)

4(X+3)-3=2(4-3X)-4 =====>>>> 4X+12-3=8-6X-4 ====>>>> 4X+9=-6X+4 ====>>>> 10X= - 5 ====>>>> X= - 1/2 ............ 4(-1/2+3)-3=2(4-3[-1/2])-4 ====>>>>>>>> 4(2 1/2) - 3 = 2(4 + 3/2) - 4 ====>>>> 10 - 3 = 2( 11/2 ) - 4 ====>>>> 7 = 11 - 4 ====>>>> 7 = 7 I HOPE THIS HELPS YOU!? ANY MORE QUESTIONS, FEEL FREE TO E MAIL ME. cats7051@yahoo.com YOUR MATH FRIEND, TERRY